### v3

# Synthetic Programming Elicitation and Repair for Text-to-Code in Very Low-Resource Programming Languages

Abstract
Recent advances in large language models (LLMs) for code applications have demonstrated remarkable zero-shot fluency and instruction following on challenging code related tasks ranging from test case generation to self-repair. Unsurprisingly, however, models struggle to compose syntactically valid programs in programming languages unrepresented in pre-training, referred to as very low-resource Programming Languages (VLPLs). VLPLs appear in crucial settings, including domain-specific languages for internal tools and tool-chains for legacy languages. Inspired by an HCI technique called natural program elicitation, we propose designing an intermediate language that LLMs “naturally” know how to use and which can be automatically compiled to a target VLPL. When LLMs generate code that lies outside of this intermediate language, we use compiler techniques to repair the code into programs in the intermediate language. Overall, we introduce synthetic programming elicitation and compilation (SPEAC), an approach that enables LLMs to generate syntactically valid code even for VLPLs. We empirically evaluate the performance of SPEAC in a case study and find that, compared to existing retrieval and fine-tuning baselines, SPEAC produces syntactically correct programs significantly more frequently without sacrificing semantic correctness.
1 Introduction
Large language models (LLMs) have demonstrated an exceptional ability to generate code from natural language prompts for popular programming languages, like Python and Java [11]. Unfortunately, these same language models struggle to generate code for low-resource programming languages, like many domain-specific languages (e.g., CUDA [32]). These challenges are even more pronounced for very low-resource programming languages (VLPLs) (e.g., UCLID5 [27]). Existing work has attempted to remedy this issue through prompting, constrained decoding, and fine-tuning strategies. Unfortunately, these approaches fail to capture the intricacies of real VLPLs and so success remains limited. To see why, consider the following three exemplars.
First, [33] include context-free grammars in text-to-code prompts to guide LLMs toward syntactically correct answers. This approach works well for simple languages but cannot capture most non-trivial programming languages, which are context-sensitive. Second, [1] use static analysis techniques to reject tokens that lead to syntactically incorrect output programs. This technique can go beyond context-free languages but assumes a linear programming process. Unfortunately, it is well known that the root cause of a programming error need not surface as it is written [26], necessitating backtracking and a nonlinear programming process. Third, [8] translate training data from high resource languages to low resource languages and then use this new data to fine-tune models. This approach is restricted to languages where the LLM is able to translate to the language reliably but unable to generate code from natural language. Further, this approach makes the overly restrictive assumption that the target low-resource language is general purpose: i.e., we cannot translate arbitrary Java programs to CUDA.
In this paper, we propose a text-to-code approach that is fundamentally different (and complementary) to prompting, decoding, and fine-tuning strategies. The first key idea behind our approach comes from natural programming elicitation, a kind of study that helps programming language designers understand how programmers “naturally” approach problems from a given programming domain [23, 9]. Programming language designers use the results of these studies to create languages that are aligned with the expectations of users, leading to less programming friction and more effective developers. We borrow this idea for the setting where LLMs are the “users” of programming languages. Akin to uncovering what human users find “natural” for a given domain, we uncover what LLMs find “natural.” Specifically, our first insight is to embrace LLM’s tendencies and design an intermediate language that aligns with these LLM expectations.
The second key idea behind our approach is that program analysis techniques that are overly aggressive for human users may be suitable for LLM “users.” For example, in UCLID5, all variables have to be declared and statically typed: an assignment like x = 0; would require a corresponding declaration like var x: integer;. But, if an LLM generates code that had an assignment without a corresponding declaration, instead of crashing, one could automatically “repair” the program and output the result.
We use these two ideas to define a new text-to-code approach called synthetic programming elicitation and compilation (SPEAC). Specifically, for a target VLPL $T$, we use synthetic programming elicitation to select an intermediate language $P$ (the “parent” language) and define a subset of the language $C$ (the “child” language). The language $P$ should be one that LLMs are good at generating (e.g. Python); the language $C$ should be easy to compile to the target VLPL $T$. Our approach takes $P$, $C$, and a compiler from $C$ to $T$, and produces a text-to-code pipeline for the VLPL $T$. This pipeline uses deductive techniques to automatically repair programs generated by LLMs that are in $P$ but not in $C$. When these deductive techniques are unable to fully repair a program, we insert a “hole” and ask an LLM to finish the repair, repeating as necessary.
We demonstrate the effectiveness of this idea in a case study that uses Python as the parent language $P$, UCLID5 [27] as the target language $T$ and a subset of Python as the child language $C$. UCLID5 is a language used for formal modeling and verification of state transition systems which has code examples numbering in the hundreds rather than thousands or millions. Furthermore, UCLID5 programs rely heavily on the notion of a transition system, which is not frequently found in other programming languages. As such, state-of-the-art LLMs are unable to generate any useful UCLID5 code out-of-the-box (see §5.1).
Overall, we make the following contributions: 1) We present SPEAC, a novel method for generating syntactically correct code from LLMs in very low resource programming languages; 2) We implement this method for the UCLID5 verification language; and 3) We demonstrate substantial improvement with SPEAC in syntactic correctness, producing parsable code in UCLID5 84.8% of the time compared to 9.1% and 3% by gpt-3.5-turbo fine-tuning or gpt-4-turbo in-context learning, respectively.
2 Related Work
LLMs for Code Generation. With the rich trove of stack overflow and github data available in public domain, proprietary models like GPT-4 [25] and open-source models such as code-llama [28] and deepseek-coder [16] have transformed code intelligence for developers. GPT-4 especially has achieved impressive success on datasets such as Mostly Basic Python Problems (MBPP) [3] and LiveCodeBench [18], 80% and 40% pass@1 results, respectively. As shown in deepseek-coder, support for less common programming languages (Bash, PHP, etc) improves substantially by inclusion in pretraining as compared to only fine-tuning as done in code-llama [16].
Training free approaches to low-resource PLs. In constrained decoding, incremental parsing is used to prevent hallucination by rejecting impossible prefixes without producing the full code [14, 1, 29, 21]. In the context of autoregressive LLMs that naturally produce code left to right, it remains an open problem how to best include the inductive bias of a grammar. Our approach advocates for allowing the LLM to hallucinate but iteratively repair errors similar to [31, 12, 24]. Orthogonal to our approach, retriever augmented generation has been applied for Dafny, another low-resource language for verification [22]. Additionally, this paper focuses on very low resourced languages (VLPLs) which have far fewer training examples in public domain.
LLMs for Low-resource PLs. Recent work has considered augmenting LLMs with support for low-resource programming languages [8, 10, 16]. [10] shows that, on smaller 125M parameter encoder-only models, fine-tuning on adjacent languages improves the monolingual performance coding tasks. However, it remains an open question whether autoregressive language models share this benefit and whether this fine-tuning trend scales to larger (7B to 70B) language models. Synthetic fine-tuning datasets curated and cleaned by LLMs have shown promise for programming tasks [8]. In particular, [8] tackles low-resource programming languages such as Julia and Rust, using an LLM to translate code examples from high-resource languages to the low-resource language. This process is promising for cases where the language model already has a baseline knowledge necessary to translate to the low-resource language which is not the case for VLPLs.
module TrafficLight {
var sigG, sigR, sigY: boolean;
var count, state: integer;
var pedestrian: boolean;
init {
sigG = false; sigY = false; sigR = true;
state = 0; count = 0; pedestrian = false; }
procedure step()
modifies sigG; modifies sigY; modifies sigR;
modifies count; modifies state; {
if (state == 0) {
sigG = false; sigY = false; sigR = true;
count = 0;
if (count < 60) { count = (count + 1); }
Figure 1: Task description from [19] (a) and output in UCLID5 (b).
3 Overview and Running Example
Given a natural language description of a programming task and a target programming language $T$, the text-to-code problem is to generate a correct program $t\in T$. Fig. 1 shows a real input-output pair generated by an instance of our approach targeting the UCLID5 programming language. Specifically, Fig. 1(a) shows a task extracted from [19], and Fig. 1(b) shows the output corresponding to that task using a prototype implementation of our approach powered by gpt-3.5-turbo-0125. Fig. 1(b) passes all compiler checks but has a subtle semantic mistake on line 4.
Fig. 2 shows the workflow that generates models based on task descriptions as shown in Fig. 1(b). The workflow is parameterized by an LLM, $L$ (e.g., gpt-3.5-turbo-0125); a target language, $T$ (e.g., UCLID5); a parent language, $P$ (e.g., Python); a child language, $C\subset P$ (e.g., a subset of Python); and a compiler, $f$, from $C$ to $T$ (e.g., a syntax-directed translation from the subset of Python to UCLID5). Given an input task, we create a prompt $q$ that asks the LLM $L$ to generate code, $p\in C$, which achieves $q$. The second step of the workflow is to repair $p$. If there is nothing wrong with $p$, or $p$ can be fixed using formal techniques described in §5.2 and §5.3, then repairing will generate a new, complete program $p^{\prime}$ and return $f(p^{\prime})$ (i.e., a program in the target language, like Fig. 1(b)). Frequently, however, repairing will generate a partial program containing “holes” (denoted “??”). For example, Fig. 5(b) shows the first $p$ generated for the task in Fig. 1(a) and Fig. 5(a) shows the corresponding partial program $p^{\prime}$ that was automatically generated using our formal techniques. Programs with holes cannot be compiled to the target language, so the third step of the workflow is ask the LLM to complete the partial program $p^{\prime}$, generating a new $p$. We use the template in Fig. 4(b) to generate the LLM prompt. Fig. 6(a) shows the output generated by gpt-3.5-turbo-0125 when asked to repair the partial program in Fig. 5(a). This program is still incorrect, but, it is now close enough that we can automatically repair it to a complete program without holes. Fig. 6(b) shows the relevant portion of that complete program. This final, complete program is directly translated to the output in Fig. 1(b). §5 elaborates on each of the workflow components.
To understand the subtle mistake in Fig. 1(b) one needs some understanding of UCLID5 [27, 30]. UCLID5 is a verification language used to model hardware and software systems as “state machine” like transition systems and to automatically check if the transition system does indeed satisfy a formal logic specification. UCLID5 transition systems primarily consist of a state space given by a set of variable declarations (e.g., lines 2-4 in Fig. 1(b)), an initialization block that represents a set of acceptable start states (e.g., lines 5-7 in Fig. 1(b)), and a transition relation block that defines how the transition system moves from one state to the next state (e.g., lines 8-29 in Fig. 1(b)). The var keyword is used to declare variables that are internal to the transition system in question while the input keyword is used to declare readonly variables that are outside the control of the transition system in question. Fig. 1(b) passes all compiler checks but has a subtle semantic mistake on line 4: var pedestrian: boolean; should be input pedestrian: boolean; because the presence of a pedestrian should not be controlled by the traffic light transition system. In §6 we mark this case as “mostly correct” because it is correct otherwise.
Figure 2: The SPEAC workflow. Users input $q$, a task in natural language, and $C$, a description of the intermediate language. The LLM takes these inputs and generates $p$, a program in $P$. We use formal techniques to repair $p$ and produce $p^{\prime}$, a program in $C$ that possibly contains holes. If $p^{\prime}$ does not contain holes, SPEAC applies $f$, a compiler from $C$ to the target language, $T$, and returns the result. Otherwise, SPEAC generates a new prompt, $q^{\prime}$, and repeats by asking the LLM to fill in the holes.
4 Background
In this section we provide the necessary technical background to understand our approach.
4.1 Algebraic Data Types and Abstract Syntax Trees
Algebraic data types (ADTs) are a representation of finite trees that is common in functional programming languages. We provide an informal definition of ADTs and point the interested reader to [5] for a more formal treatment.
An ADT consists of a set of constructors (node names), selectors (directed edge names), and testers (predicates). Each constructor has a fixed set of selectors associated with it (nodes with a particular name have a fixed set of outgoing edges). Each selector has an associated type (each edge can only point to a fixed set of node names). Unlike constructors and selectors, testers are not part of the finite trees that ADTs represent. Instead, testers are predicates that are used to ask questions about the finite trees. Specifically, every constructor is associated with exactly one unique tester: that tester returns true on a given tree iff the root of the tree is labeled with the corresponding constructor. Every instance of an ADT (a particular finite tree built from those node and edge names) must be acyclic.
For example, consider an ADT called Bool that represents simple Boolean expressions. Bool has four constructors: Not, And, True, and False. The constructor Not is associated with the selector arg, which in turn is associated with the type Bool; The constructor And is associated with the selectors arg1 and arg2, which are both associated with the type Bool; and the constructors True and False are associated with no selectors. An instance of Bool is Not(arg=And(arg1=True, arg2=False)). This is a finite tree whose root is Not and whose two leaves are True and False.
Abstract syntax trees (ASTs) are instances of ADTs—i.e., ASTs are concrete finite trees—that represent programs. Every programming language has a corresponding ADT that represents a superset of all possible programs in that programming language. Some instances of a languages’s ADT will not correspond to valid programs, e.g., if they do not additionally type check.
4.2 Satisfiability Modulo Theories and Weighted Maximum Satisfiability
Satisfiability Modulo Theories (SMT) [6] is a class of problems generalizing Boolean satisfiability (SAT) to first-order logics with additional background logical theories. We give an informal presentation of satisfiability modulo theories (SMT) that focuses on only the theory of ADTs and covers weighted maxmimum satisfiability (MAX-SMT). For more details, see [6].
[…]
The maximum satisfiability problem is, for a given (CNF) formula $\phi$, to determine the largest subset of $\phi$ that is sat (solvers aim to satisfy as many clauses as possible). The weighted maximum satisfiability problem (MAX-SMT) is a variation with two differences. First, some clauses can be “hard”—meaning they must be satisfied. Second, every “soft” clause (any clause that is not “hard”) has an associated weight. The problem is then to determine subset of $\phi$ that maximizes the sum of weights while being sat and containing all “hard” clauses.
5 Approach
In this section we describe the SPEAC approach. We begin with how to select $P$ and $C$, and then we describe the two formal techniques at the heart of our automated repair.
from FormalVerificationLibrary import Module
# GTCODE 2: creates an instance of an imported class
m = Module("myModule")
# GTCODE 3: uses a dunder method (in this case __str__) of an imported class
print(m)
Figure 3: Hypothetical LLM output with grounded theory codes as comments.
5.1 Synthetic Programming Elicitation
In this section, we present a three step recipe for conducting LLM programming studies—studies where LLMs are the subject. The results of these studies can be used to select $P$ and $C$. We call this process synthetic programming elicitation since it is inspired heavily by natural programming elicitation [23].
First Step: Setup. First prepare LLM subjects, select a target language, and collect or construct a set of programming tasks. In our case, our target language is UCLID5, our LLM subjects are gpt-4-turbo-2024-04-09 and gpt-3.5-turbo-0125, and our programming tasks are a set of regression tests written by UCLID5 developers with corresponding natural language descriptions (see §6).
The second part of the study setup is to create multiple kinds of specialized prompts. The first will ask for the output to be in the target language directly. Subsequent prompts will ask for the output to use an imaginary API in a popular host language, like Python or Java. For example, for UCLID5, we generated prompts that look like the following.
  1. 1.
    “Write a UCLID5 model for a system that counts the number of times the temperature exceeds a threshold [$\ldots$]”
  2. 2.
    “ FormalVerificationLibrary is a Python library for generating verification models [$\ldots$]. Use the FormalVerificationLibrary to model a system that counts the number of times the temperature exceeds a threshold [$\ldots$]”
Second Step: Execute. Now we execute every LLM subject on every prompt and collect the outputs. Each query should be executed independently.
Write [PARENT LANGUAGE] code to complete the following task.
> [TASK GOES HERE]
Reply with your code inside one unique code block
[DESCRIBE CHILD LANGUAGE]
I can definitely do that! Here is the code:
(a) Template for prompt $q$ in Fig. 2
[CODE WITH HOLES GOES HERE]
Make sure that your code completes the following task.
[LINES 3-10 OF (a)]
(b) Template for prompt $q^{\prime}$ in Fig. 2
Figure 4: SPEAC prompt templates for first generation (a) and hole filling (b).
Third Step: Analyze and Design. Finally, we analyze the collected LLM outputs and select $P$ and $C$ based on the analysis. To do this analysis, we follow a lightweight version of grounded theory [15].
The first step of our grounded theory analysis is to “code” the outputs generated by the LMs. In grounded theory parlance, “code” is akin to labelling parts of the generated outputs. For example, Fig. 3 shows a hypothetical LLM output for our running example along with grounded theory codes as comments. The second step is to group codes into concepts. Concepts are slightly more abstract than codes. For example, a concept that may emerge from Fig. 3 is the group of codes 1, 2, and 3. The corresponding concept could be “using a class from the hypothetical library.” In the third step, we group concepts into categories. For example, we may group concepts related to the object oriented programming paradigm as one category. Finally, we select a $P$ and $C$ that are consistent with the final categories we observed across multiple prompts.
In our study, we found that, when asked to write UCLID5 code without any special prompting or training, no LLM was able to produce code that parses (660 attempts: 10 samples per LLM per benchmark, 33 benchmarks, two LLMs). Worse still, the code generated by LLMs was inconsistent, with each LLM giving different outputs that resemble different programming languages at different times. When asked to write Python code that used a non-existent formal verification API, however, the LLM outputs were more consistent. Therefore, we selected Python as our parent language, $P$.
Specifically, the Python code was more consistent because LLM outputs fell into three broad categories, which we call “from-scratch,” “procedural,” and “object-oriented,” respectively. Programs in the “from-scratch” category used existing APIs (e.g., the Z3 solver API [13]) to re-implement what UCLID5 does, e.g., to manually model transition systems. This was the smallest significant category. Programs in the “procedural” category imported a class from the hypothetical API, created an instance of it, and then called methods from the class to declare variables, assert specifications and so on. Programs in the “object-oriented” category imported a class from the hypothetical API and extended it, including methods that correspond closely to parts of UCLID5 code.
After analyzing the outputs, we concluded that it would be easiest to translate Python programs from the “object-oriented” category to UCLID5. Therefore we defined a subset of Python that matches the “object-orineted” category and used that as our child language, $C$. Essentially, $C$ is an abstract class that the LLM must implement and extend (shown in Appx. A). Fig. 6(b) shows a portion of an example of a Python program in $C$ and Fig. 1(b) shows its corresponding UCLID5 program.
(Minimal) Prompting. Once $P$ and $C$ have been selected, we use minimal prompting to interface with LLMs. For example, for UCLID5, we use the prompt template in Fig 4(a) to create initial programs and the prompt template in Fig. 4(b) to fill in holes.
class TrafficLight(Module):
def types(self):
self.state_t = BitVector(2)
def locals(self):
self.state = BitVector(2)
self.count = BitVector(6)
self.pedestrian = Boolean()
self.sigG = Boolean()
self.sigY = Boolean()
self.sigR = Boolean()
def inputs(self):
self.pedestrian = Boolean()
def outputs(self):
self.sigG = Boolean()
self.sigY = Boolean()
self.sigR = Boolean()
def init(self):
self.state = BitVector(2)
self.count = BitVector(6)
self.pedestrian = Boolean()
self.sigG = Boolean()
self.sigY = Boolean()
self.sigR = Boolean()
def next(self):
if self.state == 0:
self.sigG = False
self.sigY = False
self.sigR = True
self.count = 0
if self.count < 60:
self.count += 1
else:
self.state = 1
elif self.state == 3:
self.sigG = False
self.sigY = False
self.sigR = False
if self.count < 60:
self.count += 1
else:
self.state = 0
class TrafficLight(Module):
def locals(self):
self.state = int
self.count = int
self.pedestrian = bool
def outputs(self):
self.sigG = bool
self.sigY = bool
self.sigR = bool
def init(self):
self.state = ??
self.count = ??
self.pedestrian = ??
self.sigG = ??
self.sigY = ??
self.sigR = ??
def next(self):
if self.state == 0:
self.sigG = False
self.sigY = False
self.sigR = True
self.count = 0
Figure 5: Response for task in Fig. 1(a) using gpt-3.5-turbo-0125 (a) and first repair (b).
5.2 Largest Consistent Subtree
Given a program $p$ in $P$, but not in $C$, our aim is to remove as little as possible from $p$ to bring it into the language $C$. That is, given the AST for $p$, we wish to generate the largest subtree of the AST, possibly containing holes, that is not rejected by the static checks of the language $C$. For example, the code in Fig.  5(b) contains some errors, including mismatches between variable and constant types (UCLID5 does not automatically cast constants and so the line self.count = 1 is interpreted as assigning the integer literal $1$ to the bitvector variable self.count, which is a type error).
We find the largest consistent subtree in three steps. First, we use an error tolerant parser to build an AST, $A$, for the language $P$. Second, beginning at the root of $A$, we recursively descend and prune anything that cannot exist in an AST for the language $C$, placing holes wherever they are needed. This is an aggressive program slice, similar to [2], which gives a new AST $A^{\prime}$. $A^{\prime}$ may or may not pass compiler checks, like type checking, or be semantically correct.
The third step finds the minimal number of AST nodes that need to be replaced with holes such that all static checks of the language $C$ pass, using the approach of [26]. Specifically, for every static check of the language, for every node of the AST, we generate a set of soft clauses. Let $S$ be the the union of all the generated clauses, and let $S^{\prime}$ be the largest subset of clauses that is sat (the output of a MAX-SMT solver). Any clause, $c$, that appears in $S$ but does not appear in $S^{\prime}$ represents a bug: if we add $c$ to $S^{\prime}$, then we will get an unsat set of clauses. So, for every such $c$, we replace the AST node that generated $c$ with a hole.
For example, in UCLID5, the two statements var x: bv32; and x = 0; would produce a set of clauses corresponding to the assertions that (a) x is a bitvector and that (b) x is an integer (since 0 is an integer). These assertions, together, are unsat, so MAX-SMT would return a set of clauses that does not include both (a) and (b). One solution would be to remove the bitvector type, producing the partial, but consistent, two statements var x: ??; and x = 0;.
We can bias the solver to include one or the other by adjusting the weights on the corresponding clauses. In practice, we assign clauses weights that are proportional to the depth of the corresponding node in the AST. This ensures that edits are closer to the leafs of the AST but it breaks syntactic minimality guarantees. In the case where weights are uniform and every node corresponds to exactly one clause, the MAX-SMT approach will make the minimal edit to the AST.
5.3 Model-Based Repair
[…]
Fig. 6 shows one buggy AST (Fig.6(a)) and the corresponding fix (Fig. 6(b)). For example, the variable count is declared as an integer in the repaired program because it is used as an integer in the buggy program. For repairs that cannot be automatically resolved by the MAX-SMT model, we use the LLM to generate code to replace the AST holes, as shown in Fig. 2.
def locals(self):
self.state = 0
self.count = 0
self.pedestrian = False
def outputs(self):
self.sigG = False
self.sigY = False
self.sigR = False
def init(self):
self.state = 0
self.count = 0
self.pedestrian = False
self.sigG = False
self.sigY = False
self.sigR = True
def locals(self):
self.count = int
self.pedestrian = bool
self.sigG = bool
self.sigR = bool
self.sigY = bool
self.state = int
def init(self):
self.state = 0
self.count = 0
self.pedestrian = False
self.sigG = False
self.sigY = False
self.sigR = True
Figure 6: Response for Fig. 5(a) using gpt-3.5-turbo-0125 (a) and second repair (b).
6 Evaluation
We implement a prototype of SPEAC for UCLID5, called Eudoxus, as a Python project that uses tree-sitter to parse and search Python programs and Z3 [13, 7] for all MAX-SMT queries [^1] Available at: https://github.com/FedericoAureliano/eudoxus. All MAX-SMT queries are solved locally on a 2.3 GHz Quad-Core Intel Core i7 with 32 GB of RAM. All LLM calls are made through the OpenAI Python API.
Benchmarks. We use three sets of UCLID5 benchmarks. The first set is a curated set of $33$ regression tests with handwritten natural language descriptions. These tests are designed to cover a broad range of UCLID5 features and were used for the synthetic programming elicitation study in § 5.1.
The second set is a set of $317$ regression tests without descriptions taken directly from the UCLID5 GitHub repository. These tests were used for fine-tuning our baseline model described below. This set is representative of the quantity and quality of data that a user may have available for a VLPL. For the purposes of fine-tuning, we use self-instruct to first ask the model to summarize the UCLID5 regression tests in natural language and then provide a natural language description using self-instruct.
The third and final set is a set of $33$ exercises and examples taken from three textbooks: [4, 20, 17]. These benchmarks cover formal modelling of concurrent systems, linear-time properties, model checking and systems with discrete dynamics. In cases where the same system is used for multiple examples, we extract the system description into a singular benchmark. We used this final set for an end-to-end evaluation.
6.1 Baselines
We compare against two LLM baselines: one-shot and fine-tuning.
One-shot. We ask the LLM to first create an outline describing each component of the UCLID5 model before lowering to code. In context, we provide an example taken from the repository that verifies a two-safety property. We demonstrate the outline process with this in-context example.
Fine-tuning. As with many VLPL, there are very limited number of NL to UCLID5 examples in the public domain, and certainly not enough for fine-tuning. We do, however, have access to $317$ regression tests from the open-source code base. We fine-tune gpt-3.5-turbo-0125 for 284 steps with learning-rate multiplier 0.16.
Semantic Score
Parse Rate 1 2 3 4 5
Eudoxus (GPT-4-turbo) 24/33 1/24 1/24 3/24 8/24 11/24
Eudoxus (GPT-3.5-turbo) 28/33 3/28 6/28 5/28 5/28 9/28
Fine-tuned GPT-3.5-turbo 2/33 0 2/2 0 0 0
One-shot with COT (GPT-4-turbo) 1/33 0 0 0 1/1 0
One-shot with COT (GPT-3.5-turbo) 0 - - - - -
One-shot (GPT-4-turbo) 0 - - - - -
One-shot (GPT-3.5-turbo) 0 - - - - -
Table 1: Eudoxus compared to fine-tuning in-context learning baselines. We report the semantic score over all correctly parsed models, where $1$ is semantically completely wrong and $5$ is semantically fully correct. Eudoxus is limited to $5$ LLM calls per benchmark, and $4$ benchmarks hit this limit. “-” indicates that this entry is not applicable for this approach.
6.2 Results
We run two variations of Eudoxus (one using gpt-3.5-turbo-0125, one using gpt-4-turbo-2024-04-09) and five variations of the baselines on all $33$ textbook benchmarks. We report the fraction of outputs that pass all compiler checks (“parse,” for short) and semantic correctness over all eight approaches in Table 1. Semantic correctness is manually assessed using a combination of manual reasoning and hand-written specifications for all programs that compile. Correctness is rated on a scale from from $1-5$, where $1$ is entirely wrong, and $4$ indicates that the model is correct with only a few minor errors (e.g., the example output described in §3 and Fig. 5). Intuitively, any score of $\geq 4$ indicates that we believe the output would be useful to text-to-code users.
Eudoxus outperforms all baselines in terms of compiler checks (see “Parse Rate” in Tab 1), passing all compiler checks on $78\%$ of benchmarks. There are four benchmarks on which Eudoxus hits the LLM call limit and fails to produce a result, and three benchmarks with small syntax errors due to bugs in our Eudoxus implementation (e.g., failing to extract code snippets from LLM outputs or printing UCLID5 code incorrectly). In contrast, we find that gpt-3.5-turbo-0125 and gpt-4-turbo-2024-04-09 are unable to produce any UCLID5 models that parse using the one-shot approach. However, gpt-4-turbo-2024-04-09 with chain-of-thought reasoning is able to produce one model that parses. The best baseline is the fine-tuned gpt-3.5-turbo-0125, but it is only able to produce two programs that parse.
For semantic correctness, we observe that $33/52$ programs that pass compiler checks produced by Eudoxus are mostly or completely semantically correct. On the other hand, only $1/3$ valid programs produced by the baselines achieves a correctness score greater than two.
7 Conclusion
We have presented a synthetic program elicitation and compilation method (SPEAC) that supports natural language to code generation for very low resource programming languages (VLPLs). The two key ideas behind SPEAC are (1) to design an interface that is “natural” for LLM “users” and (2) to use deductive techniques, which could be deemed too aggressive for human users, to automatically repair LLM outputs when possible. We implemented a prototype of SPEAC called Eudoxus that targets the UCLID5 VLPL and evaluated it on a set of $33$ benchmarks from textbooks in the same domain as UCLID5. Eudoxus performs significantly better than standard LLM baselines on syntactic correctness (from at best 2/33 to at worst 24/33) and semantic correctness (from at best 1/33 to at worst 14/33).
We believe that SPEAC can achieve similar results on other VLPLs, but further case studies are needed to evaluate that claim. In this paper, we demonstrate that our combination of synthetic programming elicitation and deductive repair can significantly improve text-to-code performance for one VLPL.
Appendix A Child Language C for UCLID5
"""An abstract class to represent a UCLID5 module."""
def types(self):
"""(Optional) Defines the type declarations.
For example, the following implementation defines a 8-bit type called T:
def types(self):
self.T = BitVector(8)
pass
def locals(self):
"""(Optional) Defines the local variables and their types.
For example, the following implementation defines an 8-bit variable x
and an integer variable y:
def locals(self):
self.x = BitVector(8)
self.y = Integer()
pass
def inputs(self):
"""(Optional) Defines the input variables and their types.
For example, the following implementation defines an input variable x,
which is an array of 8-bit bitvectors indexed by 2-bit bitvectors:
def inputs(self):
self.x = Array(BitVector(2), BitVector(8))
pass
def outputs(self):
"""(Optional) Defines the output variables and their types.
For example, the following implementation defines an output variable y,
which is a real number:
def outputs(self):
self.y = Real()
pass
def shared_vars(self):
"""(Optional) Defines the shared variables and their types.
For example, the following implementation defines a shared variable z,
which is an array of booleans indexed by integers:
def shared_vars(self):
self.z = Array(Integer(), Boolean())
pass
def instances(self):
"""(Optional) Defines the instances of other modules and relates their
input, output, and shared variables to local variables. Every instance
variable must be related to a local variable. For example, let M be
another module with inputs x and y, and output z. The following
implementation defines an instance of M called m, and connects Ms
input variable x to the local variable self.a, Ms input variable y to
the local variable self.b, and Ms output variable z to the local
variable self.c:
def instances(self):
self.m = M(x=self.a, y=self.b, z=self.c)
pass
def init(self):
"""(Optional) Defines how variables are initialized.
For example, the following implementation initializes x to 0 if y is
greater than or equal to 20:
def init(self):
if self.y >= 20:
self.x = 0
pass
def next(self):
"""(Optional) Defines the transition relation.
For example, the following implementation increments x or decrements y
depending on z:
def next(self):
if self.z:
self.x = self.x + 1
else:
self.y = self.y - 1
pass
def specification(self):
"""(Optional) Defines the specification in terms of invariant properties.
Returns:
bool: True if the specification is satisfied, False otherwise.
For example, the following implementation defines two invariants:
def specification(self):
return self.x < 10 and self.y > 0
pass
def proof(self):
"""(Optional) Defines the control block.
For example, the following implementation uses 1-induction to prove
that the specification always holds:
def proof(self):
induction(1)
pass